# Performance Analysis and Power Allocation for Massive MIMO ISAC Systems

### v2

@@ Abstract
Integrated sensing and communications (ISAC) is envisioned as a key feature in future wireless communications networks. Its integration with massive multiple-input-multiple-output (MIMO) techniques promises to leverage substantial spatial beamforming gains for both functionalities. In this work, we consider a massive MIMO-ISAC system employing a uniform planar array with zero-forcing and maximum-ratio downlink transmission schemes combined with monostatic radar-type sensing. Our focus lies on deriving closed-form expressions for the achievable communications rate and the Cramér–Rao lower bound (CRLB), which serve as performance metrics for communications and sensing operations, respectively. The expressions enable us to investigate important operational characteristics of massive MIMO-ISAC, including the mutual effects of communications and sensing as well as the advantages stemming from using a very large antenna array for each functionality. Furthermore, we devise a power allocation strategy based on successive convex approximation to maximize the communications rate while guaranteeing the CRLB constraints and transmit power budget. Extensive numerical results are presented to validate our theoretical analyses and demonstrate the efficiency of the proposed power allocation approach.
@@ I-A Related Works
Most of the aforementioned works investigated ISAC operations in MIMO scenarios [31, 12, 32, 16]. However, conventional small-sized antenna arrays may not ensure high spatial beamforming gains. In contrast, massive MIMO technology with very large arrays offers superior spectral and energy efficiency (SE/EE) for communications systems [33, 34, 35, 36]. It has been shown in [37] that by leveraging a massive MIMO radar BS, communications and radar systems can coexist with little mutual interference. Temiz et al. [38] proposed a joint uplink massive MIMO communications and orthogonal frequency-division multiplexing (OFDM) radar sensing architecture, using zero-forcing (ZF) and ordered successive interference cancellation receivers to eliminate the inter-user and radar interference during communications symbol detection. In [39, 40, 41, 42], the radar performance of massive MIMO systems is optimized under constraints on communications performance. On the other hand, the communications rate of massive MIMO-ISAC systems is maximized in [43, 30, 29] subject to constraints on the radar performance. In [44, 45], efficient channel and target parameter estimation approaches were developed for millimeter-wave massive MIMO ISAC systems, relying on compressed sampling [44] and tensor frameworks [45]. Liao et al.[46] derived closed-form expressions for the communications SINR and the sensing mainlobe-to-average-sidelobe ratio (MASR). They also proposed a power allocation algorithm to minimize the total transmit power while satisfying predefined SINR and MASR thresholds. Topal et al.[47] focused on a multi-target sensing scenario and introduced a beamforming scheme aimed at minimizing the total CRLBs for direction-of-arrival (DoA) estimates of all targets.
@@ I-B Motivation and Contributions
  • We derive closed-form expressions for the achievable rate and the CRLB for the target’s azimuth and elevation angles. These metrics are expressed as functions of the large-scale fading parameters and are unified for both the MRT and ZF schemes. The approach significantly simplifies the subsequent power allocation procedure.
  • Based on the closed-form rate and CRLB expressions, we show that the sensing objective increases the beamforming uncertainty and the inter-user interference of the communications channels for both MRT and ZF. However, these drawbacks can be mitigated by deploying a very large number of antennas. We also show that the sensing performance can be improved by increasing the number of antennas.
@@
We summarize the contributions and novelties of this paper compared to existing studies on massive MIMO ISAC systems in Table I. While SINR and achievable rate are commonly used to evaluate communications performance, the sensing performance metrics vary across the compared studies. Notably, works such as [46, 37, 38] provide closed-form expressions for the communications SINR and achievable rate. However, most of these studies evaluate sensing performance using instantaneous metrics, requiring beamforming designs that adapt to the rapid changes of small-scale fading channels. In contrast, our study is unique in deriving closed-form CRLBs for the estimation of both azimuth and elevation angles. Furthermore, although closed-form achievable rate expressions have been derived for massive MIMO communications-only systems in works such as [35, 50, 51, 52, 53, 54, 55, 56], these results are not applicable to massive MIMO ISAC systems because ISAC waveforms differ fundamentally from those used in communications-only systems.
@@
[37] Closed-form achievable rate Instantaneous probability of detection Propose using massive array at the BS to perform downlink communications and sensing/surveillance of the surrounding environment through radar scanning. Show that communications and radar scanning functions can coexist with little interference thanks to large antenna arrays.
[38] Closed-form achievable rate Instantaneous radar channel estimation mean squared error and radar image SINR Propose a dual-function radar-communications system to communicate with multiple uplink UEs while sensing targets in range using the same OFDM subcarriers. Derive expression for closed-form communications rate with perfect and imperfect CSI. Evaluate radar detection and channel estimation accuracy.
[46] Closed-form SINR Closed-form MASR Derive closed-form expressions for SINR and MASR for communications and sensing performance with MRT and ZF precoding. Calculate power allocation to minimize total transmit power under communications SINR and sensing MASR constraints.
@@
This work Closed-form achievable rate Closed-form CRLBs for target azimuth and elevation angles estimates Derive a unified closed-form expression for communications rate for both MRT and ZF precoding (not found in [47]). Derive closed-form sensing CRLBs for target azimuth and elevation angle estimates (not found in [46, 38, 37, 47]). (i) Show that sensing degrades communications rate by increasing beamforming uncertainty and inter-user interference, (ii) demonstrate how this effect can be mitigated by deploying a large number of antennas, and (iii) CRLBs approach zero when the number of antennas goes to infinity (not found in [46, 38, 37, 47]). Derive power allocation to maximize communications sum rate subject to sensing CRLB and total power constraints (different from [46, 38, 37, 47]).
@@ I-C Paper Organization and Notation
The rest of the paper is organized as follows. In Section II, we present the communications and sensing signal models and the general linear beamformers. In Section III, we derive closed-form expressions for the system communications rate and sensing CRLB, followed by important operational characteristics of the considered massive MIMO ISAC system. Section IV details the proposed power allocation method. Numerical results are given and discussed in Section V. Finally, Section VI concludes the paper.
@@ II-A1 Communications Signal Model
where ${\mathbf{W}}=[{\mathbf{w}}_{1},\ldots,{\mathbf{w}}_{K}]\in\mathbb{C}^{N_{\mathtt{t}}\times K}$ and $\bm{\Gamma}=\mathtt{diag}\left\{\sqrt{\gamma_{1}},\ldots,\sqrt{\gamma_{K}}\right\}\in\mathbb{C}^{K\times K}$ are the matrix of precoding vectors and power allocation factors, respectively, for the communications users. Specifically, ${\mathbf{w}}_{k}$ and $\gamma_{k}$ are the precoding vector and power allocated for the $k$-th communications user. On the other hand, ${\mathbf{v}}\in\mathbb{C}^{N_{\mathtt{t}}\times 1}$ and $\bar{\bm{\eta}}=[\sqrt{\eta_{1}},\ldots,\sqrt{\eta_{K}}]^{\scriptscriptstyle\mathsf{T}}\in\mathbb{C}^{K\times 1}$ are the precoding vector and power fraction allocated for sensing in each data stream. The $k$-th column of ${\mathbf{F}}$, denoted as $\mathbf{f}_{k}$, represents the dual-functional precoding vector for user $k$ and is given as $\mathbf{f}_{k}=\sqrt{\gamma_{k}}{\mathbf{w}}_{k}+\sqrt{\eta_{k}}{\mathbf{v}}$. Then, the $N_{\mathtt{t}}\times 1$ transmit signal vector during the $\ell$-th time slot is ${\mathbf{x}}_{\ell}={\mathbf{F}}{\mathbf{s}}_{\ell}=\sum_{k=1}^{K}\mathbf{f}_{k}s_{k\ell}$, and the overall dual-functional transmit waveform is denoted by ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{L}]\in\mathbb{C}^{N_{\mathtt{t}}\times L}$. Equivalently, we have ${\mathbf{X}}={\mathbf{F}}{\mathbf{S}}$.
@@ II-C Linear ISAC Beamforming
where $\tilde{{\mathbf{N}}}_{\mathtt{s}}\triangleq\alpha{\mathbf{b}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{W}}\bm{\Gamma}{\mathbf{S}}+{\mathbf{N}}_{\mathtt{s}}$. It is observed that for an arbitrary ${\mathbf{W}}\bm{\Gamma}$, the sensing beamformer ${\mathbf{v}}={\mathbf{a}}$ maximizes the received echo signal power at the BS.
@@
${\mathbf{F}}_{\mathtt{bf}}$ $={\mathbf{W}}_{\mathtt{bf}}{\boldsymbol{\Gamma}}+{\mathbf{a}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}},$ (15)
@@ III-A Communications Performance
where $\bar{\tau}\triangleq\left(\tau_{\mathtt{c}}-\tau_{\mathtt{p}}\right)/\tau_{\mathtt{c}}$. The following theorem presents the closed-form expression for the achievable rate of the communications users.
@@ III-B Sensing Performance
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\text{trace}\left(\dot{{\mathbf{G}}}_{\theta}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)\in\mathbb{R}^{1\times 1},$ (29)
$J_{\theta\phi}$ $=\kappa\left|\alpha\right|^{2}\text{trace}\left(\dot{{\mathbf{G}}}_{\phi}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)\in\mathbb{R}^{1\times 1},$ (30)
${\mathbf{J}}_{\theta\bar{\bm{\alpha}}}$ $=\kappa\mathfrak{R}{\left(\alpha^{*}\text{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)[1,j]\right)}\in\mathbb{R}^{1\times 2},$ (31)
$J_{\phi\phi}$ $=\kappa\left|\alpha\right|^{2}\text{trace}\left(\dot{{\mathbf{G}}}_{\phi}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\phi}\right)\in\mathbb{R}^{1\times 1},$ (32)
${\mathbf{J}}_{\phi\bar{\bm{\alpha}}}$ $=\kappa\mathfrak{R}{\left(\alpha^{*}\text{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\phi}\right)[1,j]\right)}\in\mathbb{R}^{1\times 2},$ (33)
${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$ $=\kappa\text{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}{\mathbf{G}}^{\scriptscriptstyle\mathsf{H}}\right)\textbf{{I}}_{2}\in\mathbb{R}^{2\times 2},$ (34)
@@
and $\kappa\triangleq\frac{2L}{\sigma_{\mathtt{s}}^{2}}$. The equivalent FIM for $(\theta,\phi)$ is
@@
${\mathbf{J}}_{\theta,\phi}=\begin{bmatrix}J_{\theta\theta}&J_{\theta\phi}\\ J_{\theta\phi}&J_{\phi\phi}-{\mathbf{J}}_{\phi\bar{\bm{\alpha}}}{\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}^{-1}{\mathbf{J}}_{\phi\bar{\bm{\alpha}}}^{\scriptscriptstyle\mathsf{T}}\end{bmatrix}=\begin{bmatrix}J_{\theta\theta}&J_{\theta\phi}\\ J_{\theta\phi}&\tilde{J}_{\phi\bar{\bm{\alpha}}}\end{bmatrix},$
@@
where $\tilde{J}_{\phi\bar{\bm{\alpha}}}\triangleq J_{\phi\phi}-{\mathbf{J}}_{\phi\bar{\bm{\alpha}}}{\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}^{-1}{\mathbf{J}}_{\phi\bar{\bm{\alpha}}}^{\scriptscriptstyle\mathsf{T}}$. The diagonal entries of ${\mathbf{J}}_{\theta,\phi}^{-1}$ serve as lower bounds on the MSE of any unbiased estimate of $\theta$ and $\phi$. As a result, we can derive the CRLBs associated with $\theta$ and $\phi$ as [61]
@@
$\mathtt{CRLB}_{\theta}\!=\!\left[\mathtt{CRLB}_{\theta,\phi}\right]_{11}=\left[{\mathbf{J}}_{\theta,\phi}^{-1}\right]_{11}=\left(J_{\theta\theta}-\frac{J_{\theta\phi}^{2}}{\tilde{J}_{\phi\bar{\bm{\alpha}}}}\right)^{-1},$ (37)
@@
$\mathtt{CRLB}_{\phi}\!=\!\left[\mathtt{CRLB}_{\theta,\phi}\right]_{22}=\left[{\mathbf{J}}_{\theta,\phi}^{-1}\right]_{22}=\left(\tilde{J}_{\phi\bar{\bm{\alpha}}}-\frac{J_{\theta\phi}^{2}}{J_{\theta\theta}}\right)^{-1}.$ (38)
We derive closed-form CRLB expressions for $\theta$ and $\phi$ with MRT and ZF precoding in the following theorem.
@@
Theorem 2
The CRLB for $\theta$ and $\phi$ in the considered massive MIMO ISAC system for either MRT or ZF precoding admits the common closed-form expression given in (39) and (40), where $\mathtt{bf}\in\{\mathtt{MRT},\mathtt{ZF}\}$, $\bm{\xi}_{\mathtt{bf}}$ is defined in (17), and
@@
See Appendix C. $\Box$
We characterize important properties of the above CRLB in the following remarks.

### v3

@@ Abstract
Integrated sensing and communications (ISAC) is envisioned as a key feature in future wireless communications networks. Its integration with massive multiple-input-multiple-output (MIMO) techniques promises to leverage substantial spatial beamforming gains for both functionalities. In this work, we consider a massive MIMO-ISAC system employing a uniform planar array with zero-forcing and maximum-ratio downlink transmission schemes combined with monostatic radar-type sensing. Our focus lies on deriving closed form expressions for the achievable communications rate and the Cramér–Rao lower bound (CRLB), which serve as performance metrics for communications and sensing operations, respectively. The expressions enable us to investigate important operational characteristics of massive MIMO-ISAC, including the mutual effects of communications and sensing as well as the advantages stemming from using a very large antenna array for each functionality. Furthermore, we devise a power allocation strategy based on successive convex approximation to maximize the communications rate while guaranteeing the CRLB constraints and transmit power budget. Extensive numerical results are presented to validate our theoretical analyses and demonstrate the efficiency of the proposed power allocation approach.
@@ I-A Related Works
Most of the aforementioned works investigated ISAC operations in MIMO scenarios [31, 12, 32, 16]. However, conventional small-sized antenna arrays may not ensure high spatial beamforming gains. In contrast, massive MIMO technology with very large arrays offers superior spectral and energy efficiency (SE/EE) for communications systems [33, 34, 35, 36]. It has been shown in [37] that by leveraging a massive MIMO radar BS, communications and radar systems can coexist with little mutual interference. Temiz et al. [38] proposed a joint uplink massive MIMO communications and orthogonal frequency-division multiplexing (OFDM) radar sensing architecture, using zero-forcing (ZF) and ordered successive interference cancellation receivers to eliminate the inter-user and radar interference during communications symbol detection. In [39, 40, 41, 42], the radar performance of massive MIMO systems is optimized under constraints on communications performance. On the other hand, the communications rate of massive MIMO-ISAC systems is maximized in [43, 30, 29] subject to constraints on the radar performance. In [44, 45], efficient channel and target parameter estimation approaches were developed for millimeter-wave massive MIMO ISAC systems, relying on compressed sampling [44] and tensor frameworks [45]. Liao et al.[46] derived closed form expressions for the communications SINR and the sensing mainlobe-to-average-sidelobe ratio (MASR). They also proposed a power allocation algorithm to minimize the total transmit power while satisfying predefined SINR and MASR thresholds. Topal et al.[47] focused on a multi-target sensing scenario and introduced a beamforming scheme aimed at minimizing the total CRLBs for direction-of-arrival (DoA) estimates of all targets.
@@ I-B Motivation and Contributions
  • We derive closed form expressions for the achievable rate and the CRLB for the target’s azimuth and elevation angles. These metrics are expressed as functions of the large-scale fading parameters and are unified for both the MRT and ZF schemes. The approach significantly simplifies the subsequent power allocation procedure.
  • Based on the closed form rate and CRLB expressions, we show that the sensing objective increases the beamforming uncertainty and the inter-user interference of the communications channels for both MRT and ZF. However, these drawbacks can be mitigated by deploying a very large number of antennas. We also show that the sensing performance can be improved by increasing the number of antennas.
@@
We summarize the contributions and novelties of this paper compared to existing studies on massive MIMO ISAC systems in Table I. While SINR and achievable rate are commonly used to evaluate communications performance, the sensing performance metrics vary across the compared studies. Notably, works such as [46, 37, 38] provide closed form expressions for the communications SINR and achievable rate. However, most of these studies evaluate sensing performance using instantaneous metrics, requiring beamforming designs that adapt to the rapid changes of small-scale fading channels. In contrast, our study is unique in deriving closed form CRLBs for the estimation of both azimuth and elevation angles. Furthermore, although closed form achievable rate expressions have been derived for massive MIMO communications-only systems in works such as [35, 50, 51, 52, 53, 54, 55, 56], these results are not applicable to massive MIMO ISAC systems because ISAC waveforms differ fundamentally from those used in communications-only systems.
@@
[37] closed form achievable rate Instantaneous probability of detection Propose using massive array at the BS to perform downlink communications and sensing/surveillance of the surrounding environment through radar scanning. Show that communications and radar scanning functions can coexist with minimal interference thanks to the use of large antenna arrays.
[38] closed form achievable rate Instantaneous radar channel estimation mean squared error and radar image SINR Propose a dual-function radar-communications system to communicate with multiple uplink UEs while sensing targets in range using the same OFDM subcarriers. Derive expression for closed form communications rate with perfect and imperfect CSI. Evaluate radar detection and channel estimation accuracy.
[46] closed form SINR closed form MASR Derive closed form expressions for SINR and MASR for communications and sensing performance with MRT and ZF precoding. Calculate power allocation to minimize total transmit power under communications SINR and sensing MASR constraints.
@@
This work closed form achievable rate closed form CRLBs for target azimuth and elevation angles estimates Derive a unified closed form expression for communications rate for both MRT and ZF precoding (not found in [47]). Derive closed form sensing CRLBs for target azimuth and elevation angle estimates (not found in [46, 38, 37, 47]). (i) Show that sensing degrades communications rate by increasing beamforming uncertainty and inter-user interference, (ii) demonstrate how this effect can be mitigated by deploying a large number of antennas, and (iii) CRLBs approach zero when the number of antennas goes to infinity (not found in [46, 38, 37, 47]). Derive power allocation to maximize communications sum rate subject to sensing CRLB and total power constraints (different from [46, 38, 37, 47]).
@@ I-C Paper Organization and Notation
The rest of the paper is organized as follows. In Section II, we present the communications and sensing signal models and the general linear beamformers. In Section III, we derive closed form expressions for the system communications rate and sensing CRLB, followed by important operational characteristics of the considered massive MIMO ISAC system. Section IV details the proposed power allocation method. Numerical results are given and discussed in Section V. Finally, Section VI concludes the paper.
@@ II-A1 Communications Signal Model
where ${\mathbf{W}}=[{\mathbf{w}}_{1},\ldots,{\mathbf{w}}_{K}]\in\mathbb{C}^{N_{\mathtt{t}}\times K}$ and $\bm{\Gamma}=\mathtt{diag}\left\{\sqrt{\gamma_{1}},\ldots,\sqrt{\gamma_{K}}\right\}\in\mathbb{C}^{K\times K}$ are the matrix of precoding vectors and power allocation factors, respectively, for the communications users. Specifically, ${\mathbf{w}}_{k}$ and $\gamma_{k}$ are the precoding vector and power allocated for the $k$-th communications user. The vectors ${\mathbf{v}}\in\mathbb{C}^{N_{\mathtt{t}}\times 1}$ and $\bar{\bm{\eta}}=[\sqrt{\eta_{1}},\ldots,\sqrt{\eta_{K}}]^{\scriptscriptstyle\mathsf{T}}\in\mathbb{C}^{K\times 1}$ represent the precoding vector and power fraction allocated for sensing in each data stream, respectively. The $k$-th column of ${\mathbf{F}}$, denoted as $\mathbf{f}_{k}$, represents the dual-functional precoding vector for user $k$ and is given as $\mathbf{f}_{k}=\sqrt{\gamma_{k}}{\mathbf{w}}_{k}+\sqrt{\eta_{k}}{\mathbf{v}}$. Then, the $N_{\mathtt{t}}\times 1$ transmit signal vector during the $\ell$-th time slot is ${\mathbf{x}}_{\ell}={\mathbf{F}}{\mathbf{s}}_{\ell}=\sum_{k=1}^{K}\mathbf{f}_{k}s_{k\ell}$, and the overall dual-functional transmit waveform is denoted by ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{L}]\in\mathbb{C}^{N_{\mathtt{t}}\times L}$. Equivalently, we have ${\mathbf{X}}={\mathbf{F}}{\mathbf{S}}$.
@@ II-C Linear ISAC Beamforming
where $\tilde{{\mathbf{N}}}_{\mathtt{s}}\triangleq\alpha{\mathbf{b}}{\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{W}}\bm{\Gamma}{\mathbf{S}}+{\mathbf{N}}_{\mathtt{s}}$. For an arbitrary ${\mathbf{W}}\bm{\Gamma}$, the optimal sensing beamformer would be ${\mathbf{v}}={\mathbf{a}}(\theta,\phi)$ if the target angles ($\theta,\phi$) were known. To maintain generality in the subsequent derivations, here we assume that ${\mathbf{v}}={\mathbf{a}}_{\mathtt{v}}(\tilde{\theta},\tilde{\phi})$ for arbitrary $\tilde{\theta}$ and $\tilde{\phi}$, so that the performance can be evaluated for the more general case where these angles are unknown.
@@
${\mathbf{F}}_{\mathtt{bf}}$ $={\mathbf{W}}_{\mathtt{bf}}{\boldsymbol{\Gamma}}+{{\mathbf{v}}}\bar{\bm{\eta}}^{\scriptscriptstyle\mathsf{T}},$ (15)
@@ III-A Communications Performance
where $\bar{\tau}\triangleq\left(\tau_{\mathtt{c}}-\tau_{\mathtt{p}}\right)/\tau_{\mathtt{c}}$. The following theorem presents the closed form expression for the achievable rate of the communications users.
@@ III-B Sensing Performance
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\mathtt{trace}\left(\dot{{\mathbf{G}}}_{\theta}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)\in\mathbb{R}^{1\times 1},$ (29)
$J_{\theta\phi}$ $=\kappa\left|\alpha\right|^{2}\mathtt{trace}\left(\dot{{\mathbf{G}}}_{\phi}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)\in\mathbb{R}^{1\times 1},$ (30)
${\mathbf{j}}_{\theta\bar{\bm{\alpha}}}$ $=\kappa\mathfrak{R}{\left(\alpha^{*}\mathtt{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\theta}\right)[1,j]\right)}\in\mathbb{R}^{1\times 2},$ (31)
$J_{\phi\phi}$ $=\kappa\left|\alpha\right|^{2}\mathtt{trace}\left(\dot{{\mathbf{G}}}_{\phi}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\phi}\right)\in\mathbb{R}^{1\times 1},$ (32)
${\mathbf{j}}_{\phi\bar{\bm{\alpha}}}$ $=\kappa\mathfrak{R}{\left(\alpha^{*}\mathtt{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}\dot{{\mathbf{G}}}^{\scriptscriptstyle\mathsf{H}}_{\phi}\right)[1,j]\right)}\in\mathbb{R}^{1\times 2},$ (33)
${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$ $=\kappa\mathtt{trace}\left({\mathbf{G}}{\mathbf{R}}_{x}{\mathbf{G}}^{\scriptscriptstyle\mathsf{H}}\right)\textbf{{I}}_{2}\in\mathbb{R}^{2\times 2},$ (34)
@@
and $\kappa\triangleq\frac{2L}{\sigma_{\mathtt{s}}^{2}}$. The equivalent FIM for $(\theta,\phi)$ is [61]
@@
${\mathbf{J}}_{\theta,\phi}=\begin{bmatrix}J_{\theta\theta}&J_{\theta\phi}\\ J_{\theta\phi}&J_{\phi\phi}-{\mathbf{j}}_{\phi\bar{\bm{\alpha}}}{\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}^{-1}{\mathbf{j}}_{\phi\bar{\bm{\alpha}}}^{\scriptscriptstyle\mathsf{T}}\end{bmatrix}=\begin{bmatrix}J_{\theta\theta}&J_{\theta\phi}\\ J_{\theta\phi}&\tilde{J}_{\phi\bar{\bm{\alpha}}}\end{bmatrix},$ (37)
where
$\tilde{J}_{\phi\bar{\bm{\alpha}}}\triangleq J_{\phi\phi}-{\mathbf{j}}_{\phi\bar{\bm{\alpha}}}{\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}^{-1}{\mathbf{j}}_{\phi\bar{\bm{\alpha}}}^{\scriptscriptstyle\mathsf{T}}.$ (38)
@@
The diagonal entries of ${\mathbf{J}}_{\theta,\phi}^{-1}$ serve as lower bounds on the MSE of any unbiased estimate of $\theta$ and $\phi$. As a result, we can derive the CRLBs associated with $\theta$ and $\phi$ with MRT and ZF precoding in the following theorem.
Theorem 2
The CRLBs for $\theta$ and $\phi$ in the considered massive MIMO ISAC system for either MRT or ZF precoding admits the common closed form expressions as follows:
@@
$\widetilde{\mathtt{CRLB}}_{\mathtt{bf},\theta}(\bm{\gamma},\rho)=\left(J_{\theta\theta}-\frac{J_{\theta\phi}^{2}}{\tilde{J}_{\phi\bar{\bm{\alpha}}}}\right)^{-1},$ (39)
@@
$\widetilde{\mathtt{CRLB}}_{\mathtt{bf},\phi}(\bm{\gamma},\rho)=\left(\tilde{J}_{\phi\bar{\bm{\alpha}}}-\frac{J_{\theta\phi}^{2}}{J_{\theta\theta}}\right)^{-1},$ (40)
where
$J_{\theta\theta}$ $=\kappa\left|\alpha\right|^{2}\Big{(}\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\left\lVert\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}+N_{\mathtt{t}}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}\right)$
$\hskip 42.67912pt+\rho{\left({\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\rVert^{2}}\lVert\dot{{\mathbf{b}}}_{\theta}\rVert^{2}+N_{\mathtt{r}}{\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\theta}\right\rVert^{2}}\right)}\Big{)}.$ (41)
$J_{\phi\phi}\!$ $=\!\kappa\!\left|\alpha\right|^{2}\Big{(}\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\left\lVert\dot{{\mathbf{a}}}_{\phi}\right\rVert^{2}\!+\!N_{\mathtt{t}}\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2}\right)$
$\hskip 42.67912pt+\rho{\left({\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\rVert^{2}}\lVert\dot{{\mathbf{b}}}_{\phi}\rVert^{2}+N_{\mathtt{r}}{\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\phi}\right\rVert^{2}}\right)}\Big{)},$ (42)
$J_{\theta\phi}\!$ $=\kappa\left|\alpha\right|^{2}\Big{(}\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}\left(N_{\mathtt{r}}\dot{{\mathbf{a}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\phi}+N_{\mathtt{t}}\dot{{\mathbf{b}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\phi}\right)+$
$\hskip 42.67912pt+\rho{\left({\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\rVert^{2}}\dot{{\mathbf{b}}}_{\phi}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{b}}}_{\theta}+N_{\mathtt{r}}\dot{{\mathbf{a}}}_{\theta}^{\scriptscriptstyle\mathsf{H}}{\mathbf{v}}{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}\dot{{\mathbf{a}}}_{\phi}\right)}\Big{)},$ (43)
${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$ $=\kappa\left(\bm{\xi}_{\mathtt{bf}}^{\scriptscriptstyle\mathsf{T}}\bm{\gamma}N_{\mathtt{t}}N_{\mathtt{r}}+\rho N_{\mathtt{r}}{\left\lVert{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}{\mathbf{a}}\right\rVert^{2}}\right)\textbf{{I}}_{2},$ (44)
${\mathbf{j}}_{\phi\bar{\bm{\alpha}}}$ $=\kappa\rho N_{\mathtt{r}}\mathfrak{R}{\left({\mathbf{a}}^{\scriptscriptstyle\mathsf{H}}{{\mathbf{v}}{\mathbf{v}}^{\scriptscriptstyle\mathsf{H}}}\dot{{\mathbf{a}}}_{\phi}[1,j]\right)},$ (45)
and $\tilde{J}_{\phi\bar{\bm{\alpha}}}$ is computed based on (38). Here, $\mathtt{bf}\in\{\mathtt{MRT},\mathtt{ZF}\}$, and $\bm{\xi}_{\mathtt{bf}}$ is defined in (17).
Proof
See Appendix C. $\Box$
Fig. 2: CRLB for $\theta$ and $\phi$ when the sensing beamformer ${\mathbf{v}}$ is determined by $\tilde{\theta}=\theta+\epsilon$ and $\tilde{\phi}=\phi+\epsilon$, $N_{\mathtt{t}}=25$, $N_{\mathtt{r}}=25$, $K=8$, $L=30$.
As explained earlier, the CRLB in (39) and (40) will depend on the choice of the sensing beamformer ${\mathbf{v}}$. Fig. 2 shows the CRLB obtained from Theorem 2 when $N_{\mathtt{t}}=N_{\mathtt{r}}=25$, $K=8$, $L=30$, and ${\mathbf{v}}$ is chosen as ${\mathbf{v}}={\mathbf{a}}(\tilde{\theta},\tilde{\phi})$, where $\tilde{\theta}=\theta+\epsilon$, $\tilde{\phi}=\phi+\epsilon$, and $\epsilon\in[0^{\circ},10^{\circ}]$. As expected, the target angle estimation performance deteriorates when the sensing beamformer is not pointed directly at the target, but we see that the CRLB is not sensitive to misalignment of the sensing beam.
The CRLB expressions in (39) and (40) are complicated due to the intricate structure of $J_{\theta\theta}$, $J_{\phi\phi}$, $J_{\theta\phi}$, ${\mathbf{J}}_{\bar{\bm{\alpha}}\bar{\bm{\alpha}}}$, ${\mathbf{j}}_{\phi\bar{\bm{\alpha}}}$, and especially $\tilde{J}_{\phi\bar{\bm{\alpha}}}$, making them intractable for analysis and even more so for optimization-based system design. In the following remark, we simplify the CRLB expression for the case where $(\tilde{\theta},\tilde{\phi})=(\theta,\phi)$, i.e., when the sensing beam is perfectly aligned with the target angles. In this case, we have ${\mathbf{v}}={\mathbf{a}}$. While this assumption does not affect the closed form expression for the achievable rate in Theorem 1, it significantly simplifies the CRLB expressions, facilitating both analysis and design.
@@
Remark 3
For the case where ${\mathbf{v}}={\mathbf{a}}$, the CRLB for $\theta$ and $\phi$ in the considered massive MIMO ISAC system for either MRT or ZF precoding can be simplified to (46) and (47), where
@@
See Appendix D. $\Box$
Based on the closed form expressions in Remark 3, we characterize important properties of the CRLB in the following remarks.